# Functional inequalities on the biased and restricted cube: an induction approach

### v2

@@ 1.2 Poincaré inequality on increasing sets
Organization. This paper is organized as follows. Section 2 is devoted to our first main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions. We develop the induction-on-dimension framework and verify the required two-point inequality in this setting. Section 3 turns to our second main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk. The arguments once again follow the induction + five-point paradigm, though the analytic verification is much simple than the biased case. Finally, Section 4 presents concluding remarks and open problems. We highlight two directions for future research: (i) developing analytic proofs of weaker forms of Harper’s edge-isoperimetric inequality, and (ii) establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
@@ 2.1 Preliminaries
In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$. One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$. Such a decomposition is motivated from Fourier expansion. This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [29]. The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$. One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality. This strategy—used in our proofs of Theorems 1.4 and 1.8—systematically reduces high-dimensional functional inequalities to finite-dimensional checks and often yields sharp constants.
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$. Let
$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
and
$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$. Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by
$g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1),$
and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$. Note that if $g$ is an increasing function, then the set $A$ is also an increasing set, since it is the support of an increasing function $g$. This implies $A_{0}\subset A_{1}$.
Fact 2.2.
@@
$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[p\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+(1-p)\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$

### v3

@@ 1.3 Organization
This paper is organized as follows. Section 2 presents our first main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk. The arguments follow the induction + five-point paradigm, though the analytic verification is much simple. Section 3 turns to our second main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions. We develop the induction-on-dimension framework and verify the required two-point inequality in this setting. Section 4 explains how inequality Eq. 8 in Theorem 1.6 leads to bounds on the mean survival (first-exit) time of random walks from increasing sets (Theorem 1.7). Section 5 verifies the mean survival time of random walks from increasing subcubes (1). Finally, Section 6 highlight a direction for future work: establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
2 Proof of Theorem 1.2
2.1 Preliminaries
In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.2. In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$. One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$. Such a decomposition is motivated from Fourier expansion. This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [Ryanbook2014]. The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$. One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality. We prove both results within the standard induction-by-restrictions, widely used in combinatorics and discrete analysis (see, e.g., [alpay2025lowerboundsdyadicsquare, beltran2023sharpisoperimetricinequalitieshypercube, BIMP2025young, CTKS2025inequalities, durcik2024sharpisoperimetricinequalitieshamming, EKL2019biased, FS2007CPCKKL, Kahnkalai07, Samorodnitsky17, yu2025averagedistancelevel1fourier]).
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}$.
Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
Lemma 2.1.
For any $f:A\to\mathbb{R}$, we have
Proof.
We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. First, separate the contribution of the last coordinate from that of the first n1n-1 coordinates; this gives
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]\\ &+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\end{split}$
Next, write $x=(x^{\prime},x_{n})$ with $x^{\prime}\in\{0,1\}^{n-1}$ and condition on $x_{n}\in\{0,1\}$. Then
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\end{split}$
Now split according to whether the flipped point stays in the same fiber $A_{b}$ for $b\in\{0,1\}$ or moves across fibers; this becomes
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
where in the second equality, we observe that
$\begin{split}\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
Lemma 2.2.
For any $f:A\to\mathbb{R}$, we have
$\begin{split}{\rm Var}_{A}[f]=&\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.\end{split}$ (11)
Proof.
Note that
$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}].\end{split}$
and
Thus we have
2.2 Induction part
Our goal is to prove that for every function $f:A\to\mathbb{R}$,
$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (13)
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$
By inductive assumption, we have
Thus our goal is to show that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (15)
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$
Thus it suffices to show that
$\begin{split}&\frac{a_{1}(c-c\sqrt{1-a_{1}})}{2a}{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}(c-c\sqrt{1-a_{0}})}{2a}{\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (16)
It is equivalent to show that
Lemma 2.3.
Proof.
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (18)
and
$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\\ &+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (20)
Now we focus on the term
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &\quad+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &\quad+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (21)
where
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\\ &\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),\end{split}$
and
Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq. 17. ∎
2.3 Five-point Inequality Part
$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})\\ &-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$
and
${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$
Thus to verify Eq. 17, it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
Note that
$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$
and
$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$
Then it suffice to prove that
$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\\ &\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &\quad\quad\ +a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &\quad\quad\ -c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (25)
Note that $a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma).$ After some simple calculations, the inequality to be verified becomes
$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (26)
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0.\end{split}$ (27)
Lemma 2.4.
$B^{2}-4AC\leq 0$.
Proof.
$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$
$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$
@@ 3.1 Preliminaries
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$. Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$. Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by $g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)$ and $g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1)$ and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$. Note that if $g$ is an increasing function, then the set $A$ is also an increasing set, since it is the support of an increasing function $g$. This implies $A_{0}\subset A_{1}$.
Fact 3.1.
@@
$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right].\end{split}$
Then we decompose each sum over $V_{1}$ and $V_{0}$ to obtain
$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$